OpenStudy (anonymous):

1.The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? A.-4, 12, -28, 60, … B.-4, -12, -28, -60, … C.-4, 12, -18, 54, … D.-4, 12, -18, -54, … 2.For the function y = f(x), what is the ordered pair for the point on the graph when x = 2m + 3? A.(x, f(m)) B.(x, 2m + 3) C.(2m + 3, f(m)) D. (2m + 3, f(2m + 3)),

OpenStudy (anonymous):

3.Given the function g(x) = 4x - 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(-4). The value of g(2) is smaller than the value of g(-4). The value of g(2) is the same as the value of g(-4). The values of g(2) and g(-4) cannot be compared.

OpenStudy (anonymous):

4. The first four terms of a sequence are shown below: 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (anonymous):

@Eli_Moses

OpenStudy (anonymous):

Can anyone help me with this

OpenStudy (anonymous):

@Johnbc